OpenStudy (anonymous):

evaluate the integral [1,2] (x^2-2/x)

OpenStudy (anonymous):

x^3/3-ln[2] 8/3-ln[2]-2/3 so 2-ln[2]

OpenStudy (anonymous):

integral of x^2 = x^3/3 integral of 2/x = 2ln(x) Over the range [1,2] means value of (x^3/3 - 2ln(x)) @ 2 minus value @ 1.

OpenStudy (anonymous):

GT is right...i messed up

OpenStudy (anonymous):

i typed it wrong it's ((x/2)-(2/x))

OpenStudy (anonymous):

Same concept. Integral of x/2 = x^2/4. Rest stands same.

OpenStudy (anonymous):